OpenStudy (anonymous):

Find the exact value of cos (2pi/3)-1/2

OpenStudy (campbell_st):

\[\cos(\frac{2\pi}{3}) = -\frac{1}{2}\] so the problem is \[- \frac{1}{2} - \frac{1}{2} = -1\]

OpenStudy (anonymous):

oh okayy :D thank you